# Short-distance thermal phase structure of charged black holes in 4D Einstein–Gauss–Bonnet gravity

### v4

@@ II.3 Entropy in the quantum regime
One of the earliest attempts to understand the microscopic degrees of freedom for black hole entropy originates from the string theory approach Strominger and Vafa (1996). However, as previously mentioned, approaches to quantum gravity and string theory through microstate counting predict additional subleading perturbative or non-perturbative contributions to the original Bekenstein-Hawking term. Nevertheless, all these approaches yield only the Bekenstein-Hawking contribution for larger geometries, with the extra terms becoming significant only in the quantum regime. A more robust and fundamental approach would be to quantize the gravitational action and deduce the resulting thermodynamic behavior. However, this task is exceedingly difficult due to the mathematical complexity involved and the uncertainty about the ultimate physical assumptions underlying quantum gravity. Yet one can adopt a more pragmatic approach by considering only the quantum corrections to entropy, thereby exploring black hole thermodynamics in the quantum regime. In this context, the Jacobson framework Jacobson (1995) and its connections to thermal fluctuations Faizal et al. (2017) may serve as a motivational basis.
The perturbative contributions to the quantum corrections of black hole entropy have the following general form Ashtekar et al. (1998); Kaul and Majumdar (2000); Pourhassan et al. (2018, 2020):
$S_{\rm p}=\mathcal{A}\ln{\left(\frac{A}{4\ell_{\rm p}^{2}}\right)}+\frac{4\mathcal{B}\ell_{\rm p}^{2}}{A}+\cdots,$ (17)
where as before $A$ denotes horizon area of the black hole while $\mathcal{A}$ and $\mathcal{B}$ are some constants related to the quantum gravity scale. Now the non-perturbative corrections read as Dabholkar et al. (2013, 2015); Chatterjee and Ghosh (2020):
@@
Here, $\eta$ is a positive parameter measuring the scale of the non-perturbative contribution to the black hole’s entropy. With that, the total BH entropy of the black hole would be $S_{\rm BH}=S_{0}+S_{\rm p}+S_{\rm np}$. There is an intriguing aspect to black holes in 4D-GB theory: the entropy already includes logarithmic contributions (albeit perturbative) from classical geometry considerations, as outlined in Ref. Fernandes (2020). Taking note of this, after incorporating exponential corrections Chatterjee and Ghosh (2020), we express the total entropy of the black hole as
@@
where $S_{0}=A+\alpha\log{A}$ is the original entropy reported in Ref. Fernandes (2020). Given that $A=r_{+}^{2}$ (omitting proportionality constants), we expand Eq. (19) as follows:
$S_{\rm exp}$ $=S_{0}+\eta\exp{\left(-S_{0}\right)}$
$=\left(M+\sqrt{M^{2}-Q^{2}-\alpha}\right)^{2}+2\alpha\log{\left(M+\sqrt{M^{2}-Q^{2}-\alpha}\right)}$
$+\eta\exp{\left[-\left(M+\sqrt{M^{2}-Q^{2}-\alpha}\right)^{2}-2\alpha\log{\left(M+\sqrt{M^{2}-Q^{2}-\alpha}\right)}\right]}.$ (20)
@@
Fig. 3 illustrates the distinctive nature of the entropy curves corresponding to different values of $\alpha$ (GB parameter) and $\eta$ (the quantum correction scale). For larger sizes, all curves tend to coincide, reflecting the dominance of the Bekenstein-Hawking term. It’s important to note that our black hole system features bifurcate horizons and exhibits distinct non-extremal and extremal geometric descriptions corresponding to the cases $M>\sqrt{Q^{2}+\alpha}$ and $M=\sqrt{Q^{2}+\alpha}$, respectively. Beyond the extremal limit, the black hole singularity becomes naked, a scenario typically forbidden by the cosmic censorship conjecture Penrose (1969). Due to the non-perturbative nature of exponential corrections in the quantum regime, the plots exhibit a sudden jump near the extremal limit whenever $M=\sqrt{Q^{2}+\alpha}$. It’s noteworthy that the entropy reaches a large value at this point but does not diverge. Upon further examination of the plots, we observe that $\eta$ contributes to a larger entropy [Fig. 3(a)], whereas $\alpha$ tends to hasten the end of evaporation (for a fixed $Q$) by shifting the extremal limit $M=\sqrt{Q^{2}+\alpha}$ each time $\alpha$ is changed [Fig. 3(b)]. Additionally, from Fig. 3(b), we notice that for the $\alpha=0$ case, the jump in entropy occurs precisely at $M=Q$, a characteristic of the RN black holes in GR.
@@
As quantum corrections to the entropy become relevant near the extremal limit $M=\sqrt{Q^{2}+\alpha}$, it is suggestive of the black hole geometric scales and the applicability of quantum corrections to the entropy. The conventional perspective of Hawking evaporation assumes that the diminishing black hole size involves the entirety of its horizon radius $r_{+}$, which naturally encompasses a combination of all three parameters in our case: $M$, $Q$, and $\alpha$. This fact is well-established concerning charged black holes Hiscock and Weems (1990), where the charge-to-mass ratio evolves as the black hole continues its evaporation towards the extremal limit. However, our definition assumes a canonical ensemble framework where $M$ fluctuates while $Q$ and $\alpha$ remain constant. This implies that the entire evaporation process occurs via $M$, which governs the black hole’s geometric scales. However, it’s important to exercise caution regarding the magnitudes of all three parameters $M$, $Q$, and $\alpha$, as their values must satisfy the quantum gravity scale in relation to entropy corrections. Here, we enforce the condition that $Q$ and $\alpha$ are extremely small, providing nearly negligible contributions. Consequently, whenever $M\rightarrow\sqrt{Q^{2}+\alpha}$, i.e., the extremal limit, our black hole tends to possess a quantum description; otherwise, it behaves as a classical system for all $M>\sqrt{Q^{2}+\alpha}$. This description would place the black hole in a coexisting phase of classical and quantum descriptions corresponding to a characteristic value of $M$. As we will later observe, this particular value of $M$ corresponds to the first root of the heat capacity $C_{Q}$ (Fig. 4). With that said, the quantum corrections to the entropy via Eq. (19) naturally follow.
@@ III Black hole stability via heat capacity
The heat capacity is displayed in Fig. 4. It’s important to note that a negative heat capacity indicates an unstable thermodynamic phase, and vice versa Davies (1978). We observe from the plots that regardless of $\alpha$ and $\eta$, $C_{Q}$ is negative for this charged black hole on larger scales [Fig. 4(c)]. The scenario unfolds differently as the black hole geometry shrinks due to Hawking evaporation. $C_{Q}$ either tends to become more negative [Fig. 4(a), $\eta=0,50$, and $80$], or transitions to positive values via $C_{Q}=0$ [Fig. 4(a), $\eta=150$, and $200$] before encountering an infinite discontinuity at a characteristic mass value $M$. This distinct behavior of $C_{Q}$ arises for specific choices of $\eta$ and $\alpha$, and may be absent in other cases [see Fig. 4(b)]. We also observe that compared to the original uncorrected case ($\eta=0$), which indicates a single root for $C_{Q}$, all curves with $\eta\neq 0$ possess two zeros of $C_{Q}$. The transition of $C_{Q}$ from negative to positive values indicates that the black hole phase changes from being unstable to a stable one, reflecting a second-order phase transition in charged black holes Davies (1978). This transition closely resembles familiar thermodynamic phase transitions in non-gravitational systems, such as ferromagnetic to paramagnetic, conductor to superconductor, liquid-crystal phase transitions, etc.
The zeros of $C_{Q}$ are typically interpreted as critical points that help distinguish between positive and negative temperature solutions Hendi and Momennia (2018). In this case, the second zero of $C_{Q}$ occurs at the extremal limit where black hole evaporation ceases, corresponding to $M=\sqrt{Q^{2}+\alpha}$. The first zero of $C_{Q}$ marks the onset of the phase transitions and represents the point where our earlier discussion regarding the definition of classical and quantum geometry for the black hole, in relation to entropy corrections, becomes relevant. This endpoint related to $M=\sqrt{Q^{2}+\alpha}$ may signify the formation of a black hole remnant, and is independent of $\eta$ in our case (our horizon geometry is $\eta$-independent). However, it manifests explicit dependence on $\alpha$, which we have numerically computed and indicated in terms of certain characteristic values of $M$ for $\alpha=0.1$ and $0.2$ in Figs. 4(a) and (b), respectively. Each time $\alpha$ takes new values, the extremal limit and corresponding shift in the last root of $C_{Q}$ is observed.
@@ IV Quantum work distribution
The thermodynamic behavior of black holes suggests the existence of a possible microstate structure. Each time a black hole emits radiation quanta due to Hawking evaporation, it alters the count of its microstates. As the black hole approaches microscopic scales, possibly near the Planck scale, quantum corrections and spacetime fluctuations become dominant. These effects manifest in the geometry of the black hole and in the associated Bekenstein-Hawking thermodynamics, as demonstrated in the preceding discussion. In this context, classical thermodynamic potentials, such as entropy and free energy, are modified due to quantum corrections. These modifications have significant implications for the partition functions and the fluctuating microstates that model black hole evaporation. In this framework, quantum work serves as a valuable quantity, quantifying the relative probabilities of transitions in a quantum system in terms of its free energy. [^2] Thus, the concept of quantum work has become a central tool in understanding quantum thermodynamics, a field grounded in non-equilibrium thermodynamic principles Fei et al. (2020); Wei (2018); Salmilehto et al. (2014).
Quantum work has been applied to study fluctuating black hole states using a stochastic approach based on Langevin and Fokker-Planck equations, as well as Jarzynski equality Iso et al. (2011); Iso and Okazawa (2011); Pourhassan et al. (2021b); Soroushfar et al. (2024); Pourhassan et al. (2022a). For macroscopic black holes, the fluctuations are negligible as the geometry is effectively classical. However, on microscopic scales, quantum corrections to entropy become significant, making quantum work a crucial factor in the thermodynamic description of the black hole. More important of these quantum corrections are the ones that manifest as exponential (non-perturbative) contributions Chatterjee and Ghosh (2020). Consequently, our goal here is to investigate the effect of non-perturbative corrections on the quantum work of the black hole. To do this, we first compute the Helmholtz free energy of the black hole. From Eq. 14, we can solve for $M$ in terms of $r_{+}$ to obtain
@@
which has been plotted in Fig.5(a) for various values of $\eta$. The plots for $F$ reveal distinct features across various size ranges. High values of $F$ for larger black hole geometric scales indicate thermodynamic instability, which is not surprising given the overall character of the black hole, particularly its larger size, which resembles the RN-like behavior noted earlier via $C_{Q}$ analysis (see Fig.4). The divergence of $F$ as $r_{+}\to 0$ reflects the true singular nature of the black hole at this limit. However, one would not expect non-zero values of $F$ at this point on physical grounds, as Hawking evaporation has removed all matter from the black hole. Thus, it is essential to interpret this as a mathematical artifact arising from the geometric properties of the black hole from which thermodynamics is defined. Clearly, the singularity near $r_{+}=0$ is a critical feature, where the geometry becomes singular. Additionally, the black hole undergoes a transition through a region of thermodynamic stability, marked by negative free energy, as it evaporates from larger unstable configurations toward the divergent $F$ at $r_{+}=0$. These stable/unstable phase transitions are a common feature in charged black hole thermodynamics, resulting from the competition between charge and mass Davies (1978); Masood A. S. Bukhari et al. (2023).
Between two successive evaporations, the black hole reduces in size from an initial horizon radius $r_{i}$ to a final radius $r_{f}$. The corresponding change in free energy is given by $\Delta F=F_{\rm f}-F_{\rm i}$. Quantum work is generally defined through the expression Jarzynski (1997)
$\left\langle\exp{\left(-\frac{W}{T_{\rm BH}}\right)}\right\rangle=\exp{\left(-\frac{\Delta F}{T_{\rm BH}}\right)}.$ (24)
This expression essentially relates the average work $W$ done on the black hole system under non-equilibrium conditions to the equilibrium free energy $F$. Furthermore, the Jensen inequality $\langle e^{x}\rangle\geq e^{\langle x\rangle}$ guarantees that $\langle W\rangle-\Delta F\geq 0$, which reflects the second law of thermodynamicsIso and Okazawa (2011). In broader frameworks, quantum work has been proposed as a tool to explore information preservation in Hawking effect Pourhassan et al. (2021b).
@@ V.1 Basic tenets of Ruppeiner geometry
Information or thermodynamic geometry, or in short, geometrothermodynamics, provides tools for understanding phase transitions and the stability of systems undergoing fluctuations around thermal equilibrium. In this approach, a parameter space, akin to Riemannian geometry in gravitational systems, is defined, spanning some extensive quantities of the system, which later aids in defining a Ricci-like curvature Ruppeiner (1995). The initial impetus came from Weinhold Weinhold (1975a, b), who defined the metric by taking the Hessian of the internal energy with respect to the other extensive variables of the system. This was followed by a rigorous approach by Ruppeiner Ruppeiner (1979, 1981, 2014), who employed entropy instead of internal energy. While these methods have been primarily developed and rigorously applied in various well-known fluctuating systems such as quantum liquids, magnetic systems, Ising models, and so on Ruppeiner (1979, 1981); Janyszek and Mrugala (1989); Ruppeiner (2014), their applicability however spans more exotic systems, such as black holes Ruppeiner (2014). It is anticipated that such a construction may potentially offer insights into the microscopic structure of black hole thermodynamics, which is typically absent in the conventional Bekenstein-Hawking formalism.
@@
Denoting the internal energy by $M$, Weinhold metric has the form $g_{\mu\nu}^{W}=\partial_{\mu}\partial_{\nu}M(S,N^{i}),$ where $S$ is the entropy, while $N^{i}$ are all other extensive quantities indexed by $i$. These quantities may include volume, internal energy etc. Each combination of $\mu,\nu=0,1,2,\dots$ represents one of these quantities. With this, Weinhold line element is given by $\mathrm{d}s_{W}^{2}=g_{\mu\nu}^{W}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}.$ Likewise, we have Ruppeiner metric given by $g_{\mu\nu}=-\frac{\partial^{2}S}{\partial x^{\mu}\partial x^{\nu}}.$ Several other information geometric methods have been developed recently, such as Quevedo Quevedo (2007); Quevedo and Sanchez (2008) and the Hendi-Panahiyan-Eslam-Panah-Momennia (HPEM) metric Hendi et al. (2015), which have also proven useful for exploring black hole thermodynamics. However, our focus here is on employing the Ruppeiner formalism, for which we will now provide a detailed derivation for context and clarity.
Let’s start with the standard Boltzmann entropy formula $S=k_{\rm B}\ln\Omega$, with $\Omega$ being the microstate count $\Omega=\exp{\left(S/k_{\rm B}\right)}.$ Next, we consider a parameter space comprising $x^{0}$ and $x^{1}$ that define the black hole. As fluctuations occur in the system, we can estimate the probability of finding the black hole system within the intervals $x^{0}+\mathrm{d}x^{0}$ and $x^{1}+\mathrm{d}x^{1}$ as $P(x^{0},x^{1})\mathrm{d}x^{0}\mathrm{d}x^{1}=\Lambda\Omega(x^{0},x^{1})\mathrm{d}x^{0}\mathrm{d}x^{1}$, where $\Lambda$ is a normalization constant. Using the expression for $\Omega$ given above, we may write $P(x^{0},x^{1})\propto\exp{\left(S/k_{\rm B}\right)}$ and $S(x^{0},x^{1})=S_{\rm BH}(x^{0},x^{1})+S_{\rm E}(x^{0},x^{1}),$ with $S_{\rm BH}$ and $S_{\rm E}$ respectively being the black hole and environment entropies. If there is a small fluctuation in the equilibrium entropy around the point $x_{0}^{\mu}$ (with $\mu,\nu=0,1$), one can Taylor expand the total entropy around $x_{0}^{\mu}$ as follows:
$S$ $=S_{0}+\frac{\partial S_{\rm BH}}{\partial x^{\mu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm bh}^{\mu}+\frac{\partial S_{\rm E}}{\partial x^{\mu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm E}^{\mu}$
$+\frac{1}{2}\frac{\partial^{2}S_{\rm BH}}{\partial x^{\mu}\partial x^{\nu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm bh}^{\mu}\Delta x_{\rm bh}^{\nu}+\frac{1}{2}\frac{\partial^{2}S_{\rm E}}{\partial x^{\mu}\partial x^{\nu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm E}^{\mu}\Delta x_{\rm E}^{\nu}+\cdots,$ (25)
@@
$\Delta S$ $=\frac{1}{2}\frac{\partial^{2}S_{\rm BH}}{\partial x^{\mu}\partial x^{\nu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm bh}^{\mu}\Delta x_{\rm bh}^{\nu}+\frac{1}{2}\frac{\partial^{2}S_{\rm E}}{\partial x^{\mu}\partial x^{\nu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm E}^{\mu}\Delta x_{\rm E}^{\nu}.$ (27)
As the entropy of the environment is almost equal to the total entropy, i.e., $S_{\rm E}\sim S$, the corresponding fluctuations in $S_{\rm E}$ are negligible. Consequently, we are left with only the black hole system, such that $P(x^{0},x^{1})\propto\exp{\left(-\frac{1}{2}\Delta l^{2}\right)},$ with $\Delta l^{2}$ given by $\Delta l^{2}=-\frac{1}{k_{\rm B}}g_{\mu\nu}\Delta x^{\mu}\Delta x^{\nu}.$ Setting $k_{\rm B}=1$, one has $\Delta l^{2}=g_{\mu\nu}\Delta x^{\mu}\Delta x^{\nu}$, where $g_{\mu\nu}=-\frac{\partial^{2}}{\partial x^{\mu}\partial x^{\nu}}S_{\rm BH}.$ Given that probability is a dimensionless scalar, $\Delta l^{2}$, as given up, is a dimensionless, positive definite, invariant quantity. This line element mimics the line element in black holes and is usually considered a quantifying measure of the thermodynamic length between two fluctuating black hole microstates. Quoting Ruppeiner Ruppeiner (2014): “Thermodynamic states are further apart if the fluctuation probability is less.” This principle resonates with Le Chatelier’s principle, which ensures the local stability of thermodynamic systems. Dropping the subscript $\rm BH$, we write $g_{\mu\nu}=-\frac{\partial^{2}}{\partial x^{\mu}\partial x^{\nu}}S$, as the metric of Ruppeiner geometry. Based on this metric, we can now compute the associated curvature scalar in the same fashion as one usually does in Riemannian geometry. Given the Christoffel connections are $\Gamma_{\mu\nu}^{\sigma}=\frac{1}{2}g^{\sigma\rho}\left(\partial_{\nu}g_{\rho\mu}+\partial_{\mu}g_{\rho\nu}-\partial_{\rho}g_{\mu\nu}\right)$, along with the Riemann tensor $R_{\rho\mu\nu}^{\sigma}=\partial_{\nu}\Gamma_{\rho\mu}^{\sigma}-\partial_{\mu}\Gamma_{\rho\nu}^{\sigma}+\Gamma_{\rho\mu}^{\delta}\Gamma_{\delta\nu}^{\sigma}-\Gamma_{\rho\nu}^{\delta}\Gamma_{\delta\mu}^{\sigma}$, one can define Ricci tensor and scalar as $R_{\mu\nu}=R_{\mu\sigma\nu}^{\sigma}$ and $R=g^{\mu\nu}R_{\mu\nu}$. Here, the Ricci curvature is Carroll (2019)

### v5

@@ II.3 Entropy in the quantum regime
One of the earliest attempts to understand the microscopic degrees of freedom for black hole entropy originates from the string theory approach [113]. However, as previously mentioned, approaches to quantum gravity and string theory through microstate counting predict additional subleading perturbative or non-perturbative contributions to the original Bekenstein-Hawking term. Nevertheless, all these approaches yield only the Bekenstein-Hawking contribution for larger geometries, with the extra terms becoming significant only in the quantum regime. A more robust and fundamental approach would be to quantize the gravitational action and deduce the resulting thermodynamic behavior. However, this task is exceedingly difficult due to the mathematical complexity involved and the uncertainty about the ultimate physical assumptions underlying quantum gravity. Yet one can adopt a more pragmatic approach by considering only the quantum corrections to entropy, thereby exploring black hole thermodynamics in the quantum regime. In this context, the Jacobson framework [63] and its connections to thermal fluctuations [32] may serve as a motivational basis.
The perturbative contributions to the quantum corrections of black hole entropy have the following general form [10, 66, 90, 94]:
[…]
where as before $A$ denotes horizon area of the black hole while $\mathcal{A}$ and $\mathcal{B}$ are some constants related to the quantum gravity scale. Now the non-perturbative corrections read as [27, 28, 25]:
@@
Here, $\eta$ is a positive parameter measuring the scale of the non-perturbative contribution to the black hole’s entropy. With that, the total BH entropy of the black hole would be $S_{\rm BH}=S_{0}+S_{\rm p}+S_{\rm np}$. There is an intriguing aspect to black holes in 4D-GB theory: the entropy already includes logarithmic contributions (albeit perturbative) from classical geometry considerations, as outlined in Ref. [35]. Taking note of this, after incorporating exponential corrections [25], we express the total entropy of the black hole as
@@
[…]
@@
Fig. 3 illustrates the distinctive nature of the entropy curves corresponding to different values of $\alpha$ (GB parameter) and $\eta$ (the quantum correction scale). For larger sizes, all curves tend to coincide, reflecting the dominance of the Bekenstein-Hawking term. It’s important to note that our black hole system features bifurcate horizons and exhibits distinct non-extremal and extremal geometric descriptions corresponding to the cases $M>\sqrt{Q^{2}+\alpha}$ and $M=\sqrt{Q^{2}+\alpha}$, respectively. Beyond the extremal limit, the black hole singularity becomes naked, a scenario typically forbidden by the cosmic censorship conjecture [89]. Due to the non-perturbative nature of exponential corrections in the quantum regime, the plots exhibit a sudden jump near the extremal limit whenever $M=\sqrt{Q^{2}+\alpha}$. It’s noteworthy that the entropy reaches a large value at this point but does not diverge. Upon further examination of the plots, we observe that $\eta$ contributes to a larger entropy [Fig. 3(a)], whereas $\alpha$ tends to hasten the end of evaporation (for a fixed $Q$) by shifting the extremal limit $M=\sqrt{Q^{2}+\alpha}$ each time $\alpha$ is changed [Fig. 3(b)]. Additionally, from Fig. 3(b), we notice that for the $\alpha=0$ case, the jump in entropy occurs precisely at $M=Q$, a characteristic of the RN black holes in GR.
@@
As quantum corrections to the entropy become relevant near the extremal limit $M=\sqrt{Q^{2}+\alpha}$, it is suggestive of the black hole geometric scales and the applicability of quantum corrections to the entropy. The conventional perspective of Hawking evaporation assumes that the diminishing black hole size involves the entirety of its horizon radius $r_{+}$, which naturally encompasses a combination of all three parameters in our case: $M$, $Q$, and $\alpha$. This fact is well-established concerning charged black holes [57], where the charge-to-mass ratio evolves as the black hole continues its evaporation towards the extremal limit. However, our definition assumes a canonical ensemble framework where $M$ fluctuates while $Q$ and $\alpha$ remain constant. This implies that the entire evaporation process occurs via $M$, which governs the black hole’s geometric scales. However, it’s important to exercise caution regarding the magnitudes of all three parameters $M$, $Q$, and $\alpha$, as their values must satisfy the quantum gravity scale in relation to entropy corrections. Here, we enforce the condition that $Q$ and $\alpha$ are extremely small, providing nearly negligible contributions. Consequently, whenever $M\rightarrow\sqrt{Q^{2}+\alpha}$, i.e., the extremal limit, our black hole tends to possess a quantum description; otherwise, it behaves as a classical system for all $M>\sqrt{Q^{2}+\alpha}$. This description would place the black hole in a coexisting phase of classical and quantum descriptions corresponding to a characteristic value of $M$. As we will later observe, this particular value of $M$ corresponds to the first root of the heat capacity $C_{Q}$ (Fig. 4). With that said, the quantum corrections to the entropy via Eq. (19) naturally follow.
@@ III Black hole stability via heat capacity
The heat capacity is displayed in Fig. 4. It’s important to note that a negative heat capacity indicates an unstable thermodynamic phase, and vice versa [30]. We observe from the plots that regardless of $\alpha$ and $\eta$, $C_{Q}$ is negative for this charged black hole on larger scales [Fig. 4(c)]. The scenario unfolds differently as the black hole geometry shrinks due to Hawking evaporation. $C_{Q}$ either tends to become more negative [Fig. 4(a), $\eta=0,50$, and $80$], or transitions to positive values via $C_{Q}=0$ [Fig. 4(a), $\eta=150$, and $200$] before encountering an infinite discontinuity at a characteristic mass value $M$. This distinct behavior of $C_{Q}$ arises for specific choices of $\eta$ and $\alpha$, and may be absent in other cases [see Fig. 4(b)]. We also observe that compared to the original uncorrected case ($\eta=0$), which indicates a single root for $C_{Q}$, all curves with $\eta\neq 0$ possess two zeros of $C_{Q}$. The transition of $C_{Q}$ from negative to positive values indicates that the black hole phase changes from being unstable to a stable one, reflecting a second-order phase transition in charged black holes [30]. This transition closely resembles familiar thermodynamic phase transitions in non-gravitational systems, such as ferromagnetic to paramagnetic, conductor to superconductor, liquid-crystal phase transitions, etc.
The zeros of $C_{Q}$ are typically interpreted as critical points that help distinguish between positive and negative temperature solutions [54]. In this case, the second zero of $C_{Q}$ occurs at the extremal limit where black hole evaporation ceases, corresponding to $M=\sqrt{Q^{2}+\alpha}$. The first zero of $C_{Q}$ marks the onset of the phase transitions and represents the point where our earlier discussion regarding the definition of classical and quantum geometry for the black hole, in relation to entropy corrections, becomes relevant. This endpoint related to $M=\sqrt{Q^{2}+\alpha}$ may signify the formation of a black hole remnant, and is independent of $\eta$ in our case (our horizon geometry is $\eta$-independent). However, it manifests explicit dependence on $\alpha$, which we have numerically computed and indicated in terms of certain characteristic values of $M$ for $\alpha=0.1$ and $0.2$ in Figs. 4(a) and (b), respectively. Each time $\alpha$ takes new values, the extremal limit and corresponding shift in the last root of $C_{Q}$ is observed.
@@ IV Quantum work distribution
The thermodynamic behavior of black holes suggests the existence of a possible microstate structure. Each time a black hole emits radiation quanta due to Hawking evaporation, it alters the count of its microstates. As the black hole approaches microscopic scales, possibly near the Planck scale, quantum corrections and spacetime fluctuations become dominant. These effects manifest in the geometry of the black hole and in the associated Bekenstein-Hawking thermodynamics, as demonstrated in the preceding discussion. In this context, classical thermodynamic potentials, such as entropy and free energy, are modified due to quantum corrections. These modifications have significant implications for the partition functions and the fluctuating microstates that model black hole evaporation. In this framework, quantum work serves as a valuable quantity, quantifying the relative probabilities of transitions in a quantum system in terms of its free energy. [^2] Thus, the concept of quantum work has become a central tool in understanding quantum thermodynamics, a field grounded in non-equilibrium thermodynamic principles [33, 118, 107].
Quantum work has been applied to study fluctuating black hole states using a stochastic approach based on Langevin and Fokker-Planck equations, as well as Jarzynski equality [61, 62, 97, 112, 91]. For macroscopic black holes, the fluctuations are negligible as the geometry is effectively classical. However, on microscopic scales, quantum corrections to entropy become significant, making quantum work a crucial factor in the thermodynamic description of the black hole. More important of these quantum corrections are the ones that manifest as exponential (non-perturbative) contributions [25]. Consequently, our goal here is to investigate the effect of non-perturbative corrections on the quantum work of the black hole. To do this, we first compute the Helmholtz free energy of the black hole. From Eq. 14, we can solve for $M$ in terms of $r_{+}$ to obtain
@@
which has been plotted in Fig.5(a) for various values of $\eta$. The plots for $F$ reveal distinct features across various size ranges. High values of $F$ for larger black hole geometric scales indicate thermodynamic instability, which is not surprising given the overall character of the black hole, particularly its larger size, which resembles the RN-like behavior noted earlier via $C_{Q}$ analysis (see Fig.4). The divergence of $F$ as $r_{+}\to 0$ reflects the true singular nature of the black hole at this limit. However, one would not expect non-zero values of $F$ at this point on physical grounds, as Hawking evaporation has removed all matter from the black hole. Thus, it is essential to interpret this as a mathematical artifact arising from the geometric properties of the black hole from which thermodynamics is defined. Clearly, the singularity near $r_{+}=0$ is a critical feature, where the geometry becomes singular. Additionally, the black hole undergoes a transition through a region of thermodynamic stability, marked by negative free energy, as it evaporates from larger unstable configurations toward the divergent $F$ at $r_{+}=0$. These stable/unstable phase transitions are a common feature in charged black hole thermodynamics, resulting from the competition between charge and mass [30, 85].
Between two successive evaporations, the black hole reduces in size from an initial horizon radius $r_{i}$ to a final radius $r_{f}$. The corresponding change in free energy is given by $\Delta F=F_{\rm f}-F_{\rm i}$. Quantum work is generally defined through the expression [65]
[…]
This expression essentially relates the average work $W$ done on the black hole system under non-equilibrium conditions to the equilibrium free energy $F$. Furthermore, the Jensen inequality $\langle e^{x}\rangle\geq e^{\langle x\rangle}$ guarantees that $\langle W\rangle-\Delta F\geq 0$, which reflects the second law of thermodynamics[62]. In broader frameworks, quantum work has been proposed as a tool to explore information preservation in Hawking effect [97].
@@ V.1 Basic tenets of Ruppeiner geometry
Information or thermodynamic geometry, or in short, geometrothermodynamics, provides tools for understanding phase transitions and the stability of systems undergoing fluctuations around thermal equilibrium. In this approach, a parameter space, akin to Riemannian geometry in gravitational systems, is defined, spanning some extensive quantities of the system, which later aids in defining a Ricci-like curvature [105]. The initial impetus came from Weinhold [122, 121], who defined the metric by taking the Hessian of the internal energy with respect to the other extensive variables of the system. This was followed by a rigorous approach by Ruppeiner [103, 104, 106], who employed entropy instead of internal energy. While these methods have been primarily developed and rigorously applied in various well-known fluctuating systems such as quantum liquids, magnetic systems, Ising models, and so on [103, 104, 64, 106], their applicability however spans more exotic systems, such as black holes [106]. It is anticipated that such a construction may potentially offer insights into the microscopic structure of black hole thermodynamics, which is typically absent in the conventional Bekenstein-Hawking formalism.
@@
Denoting the internal energy by $M$, Weinhold metric has the form $g_{\mu\nu}^{W}=\partial_{\mu}\partial_{\nu}M(S,N^{i}),$ where $S$ is the entropy, while $N^{i}$ are all other extensive quantities indexed by $i$. These quantities may include volume, internal energy etc. Each combination of $\mu,\nu=0,1,2,\dots$ represents one of these quantities. With this, Weinhold line element is given by $\mathrm{d}s_{W}^{2}=g_{\mu\nu}^{W}\mathrm{d}x^{\mu}\mathrm{d}x^{\nu}.$ Likewise, we have Ruppeiner metric given by $g_{\mu\nu}=-\frac{\partial^{2}S}{\partial x^{\mu}\partial x^{\nu}}.$ Several other information geometric methods have been developed recently, such as Quevedo [99, 98] and the Hendi-Panahiyan-Eslam-Panah-Momennia (HPEM) metric [55], which have also proven useful for exploring black hole thermodynamics. However, our focus here is on employing the Ruppeiner formalism, for which we will now provide a detailed derivation for context and clarity.
[…]
$S=S_{0}+\frac{\partial S_{\rm BH}}{\partial x^{\mu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm bh}^{\mu}+\frac{\partial S_{\rm E}}{\partial x^{\mu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm E}^{\mu}\\ +\frac{1}{2}\frac{\partial^{2}S_{\rm BH}}{\partial x^{\mu}\partial x^{\nu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm bh}^{\mu}\Delta x_{\rm bh}^{\nu}\\ +\frac{1}{2}\frac{\partial^{2}S_{\rm E}}{\partial x^{\mu}\partial x^{\nu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm E}^{\mu}\Delta x_{\rm E}^{\nu}+\cdots,$ (25)
@@
$\Delta S=\frac{1}{2}\frac{\partial^{2}S_{\rm BH}}{\partial x^{\mu}\partial x^{\nu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm bh}^{\mu}\Delta x_{\rm bh}^{\nu}\\ +\frac{1}{2}\frac{\partial^{2}S_{\rm E}}{\partial x^{\mu}\partial x^{\nu}}\bigg|_{x^{\mu}=x_{0}^{\mu}}\Delta x_{\rm E}^{\mu}\Delta x_{\rm E}^{\nu}.$ (27)
[…]